OpenStudy (anonymous):

write an equation in slope-intercept form of a line containing the points (-8,9) and (-6,-5) a) y=x-11 b)y=-7x+37 c)y=-7-47 d)y=-x-1

OpenStudy (anonymous):

m=14/-2=-7 y=-7x+7(-6)-5=-7x-47

OpenStudy (anonymous):

\[\frac{y_2-y_1}{x_2-x_1}=\frac{9-(-5)}{-8-(-6)}=\frac{9+5}{-8+6}=\frac{14}{-2}=-7\]

OpenStudy (anonymous):

so the answer s not in the selection.. to get the slope m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

i guess x is missing in option c

OpenStudy (anonymous):

\[y-9=-7(x-(-8))\] \[y-9=-7(x+8)\] \[y-9=-7x-56\] \[y-9=-7x-49\]

OpenStudy (anonymous):

oops \[y=-7x-47\] uzma way snappier!